OpenStudy (linn99123):

Which is a solution to the equation 6x + y = –4? A. (1, –10) B. (2, 8) C. (1, 2) D. (2, –8)

OpenStudy (anonymous):

6x + y = -4; solve for y y = -6x - 4: plug and chug y = -6(2) - 4 y = -12 - 4 y = -16...nope y = -6(1) - 4 y = -6 - 4 y = -10..yes!

OpenStudy (linn99123):

so A>?

OpenStudy (anonymous):

yea :)